An aqueous solution contains 0.352 M hydrocyanic acid.

Calculate the pH of the solution after the addition of 4.45E-2 moles of potassium hydroxide to 250 mL of this solution.
(Assume that the volume change does not change upon adding potassium hydroxide)
Answer #1

HCN + KOH ---------------> KCN + H2O

[HCN] = 0.352M,

[KOH] = 4.45 x 10-2 x 1000/250 = 17.8 x 10-2 = 0.178M

Acid concentration > base concentration

[H+]= 0.352 - 0.178 /2 = 0.174/2M = 0.087M

PH = -log[0.087] = 1.06

PH = 1.06